Cinco De Mayo has a target debt-equity ratio of 0.69. (Hint: if you're not sure what to do about "debt-equity ratio", watch my YouTube video on that topic!) The yield to maturity on its bonds is 11 percent. Its cost of equity is 17 percent. The corporate income tax rate is 34 percent. Calculate the WACC for this company.


rev: 09_20_2012

Multiple Choice

  • 11.24%

  • 10.28%

  • 12.37%

  • 13.02%

  • 13.67%

This is the last question in the assignment. To submit, use Alt + S. To access other questions, proceed to the question map button.Next Visit question map

Question10of10Total10 of 10

Prev
Answer #1

WACC = weight of debt * after tax cost of debt + weight of equity * cost of equity

debt/equity = 0.69

debt + equity = 1

=>

debt = 69/169

equity = 100/169

WACC = 69/169 * 11% * (1-0.34) + 100/169 * 17%

= 13.02%